# Non-Fermi liquid behaviour of CDW instabilities in fractionally-filled moiré flatbands

### v1

@@ C.2 Computation of $\gamma_{d-1}$-dependent part
The leading order dependence of $\Sigma_{2}(k)$ on $k_{d}$ and $k_{d-1}$ can be extracted by setting $\mathbf{K}=0$. Hence, we will evaluate
@@
where $\delta_{q-k}=q_{d-1}-k_{d-1}+k_{d}^{2}+q_{d}^{2}-2\,k_{q}\,q_{d}$. Here,
$I_{3}$ $\simeq\int_{q}\frac{\delta_{q-k}}{\mathbf{Q}^{2}+\delta^{2}_{q-k}}\times\frac{1}{q_{d}^{2}}=\int\frac{d^{d-1}\mathbf{Q}\,dq_{d}\,du}{(2\,\pi)^{d+1}}\frac{u-k_{d-1}}{2\,\left[|\mathbf{Q}|^{2}+\left(u-k_{d-1}\right)^{2}\right]}\times\frac{2}{q_{d}^{2}}\quad[\text{where }u=\delta_{q}+2\,k_{d}\,q_{d}+k_{d}^{2}]\,.$ (52)
Using the identity
[…]
for the Cauchy principal value of the integral, the $q_{d}$-integral is performed first to obtain the form
@@
$=\int\frac{d^{d-1}\mathbf{Q}\,d\tilde{u}}{(2\,\pi)^{d}}\frac{k_{d}^{2}+\tilde{u}-k_{d-1}}{\left[|\mathbf{Q}|^{2}+\left(k_{d}^{2}+\tilde{u}-k_{d-1}\right)^{2}\right]}\times\frac{\Theta(\tilde{u})}{2\,\sqrt{\tilde{u}}}\quad[\text{where }\tilde{u}=u-k_{d}^{2}]$
@@
The factor $\Gamma\big(\frac{3}{2}-d\big)$ has a pole at $d=3/2$, which shows that this term has (1) a logarithmic divergence at $d=3/2$, and (2) a linear divergence at $d=5/2$, when we translate the divergences in the language of the Wilsonian cutoff $\Lambda\sim\mu$. We need to treat this result carefully by remembering that, in the dimensional regularization procedure, UV divergences of all degrees show up as the poles of $\Gamma$-functions. The degree of divergence can be understood by using an explicit UV cut-off in the language of the Wilsonian RG, which is denoted here by $\Lambda$. Although this term will play an important role for analyzing UV-stable fixed points, it should be discarded here because we are considering the RG flows in the IR, and this term represents an IR-irrelevant operator for the theory in $d=5/2-\epsilon$.
@@ Appendix D Renormalization Group flows under minimal subtraction scheme
Eq. (II) is supposed to be the physical action, defined at an energy scale $\mu\sim\Lambda$, consisting of the fundamental Lagrangian with non-divergent quantities. However, we have seen that the loop integrals lead to divergent terms and, in order to cure it, we employ the renormalization procedure, using dimensional regularization as the regularization method. In our dimensional regularization formalism, the UV-divergent terms are the ones arising in the $\epsilon\rightarrow 0$ limit. We use the minimal subtraction ($\rm MS$}) renormalization scheme to control the UV divergences [52, 53], which involves cancelling the divergent parts of the loop-contributions via adding appropriate counterterms. More precisely, we adopt the modified minimal subtraction ($\overline{\rm MS}$) scheme where, in addition to the divergent term, we absorb the universal term proportional to $\epsilon^{0}$ (that always accompanies the term with the $1/\epsilon$ pole) into the corresponding counterterm.
The action, consisting of the counterterms to absorb the singular terms, takes the form:
@@
The counterterm-factors are given by the power series
@@
Subtracting $\mathcal{S}_{CT}$ from the so-called bare action $S_{\text{bare}}$, we obtain the renormalized action, which is the physical effective action of the theory, re-written in terms of non-divergent quantum parameters. While the bare parameters can be divergent, the physical observables are the renormalized coupling constants, which are determined by the RG equations. The RG flows describe the evolution of the bare couplings as functions of the floating energy scale $\mu\,g^{-l}$ (i.e., with respect to an increasing logarithmic length scale $l$). To achieve this objective, we first define the bare (or fundamental) action
@@
consisting of the bare quantities, where the superscript “BB” has been used to denote the bare fields, couplings, frequency, and momenta. We now relate the bare quantities to the so-called renormalized quantities (without the superscript “BB”) via the multiplicative $Z_{\zeta}$-factors such that
@@
Observing that there exists a freedom to change the renormalization of the fields and the renormalization of momenta without affecting the action, we have exploited it by requiring $k^{B}_{d}=k_{d}$, which is equivalent to measuring the scaling dimensions of all the other quantities relative to the scaling dimension of $k_{d}$. $\mathcal{S}$ now represents the renormalized action (also known as the Wilsonian effective action) because it consists of the renormalized quantities. Basically, we have written the fundamental action of our theory in two different ways [54], which allows us to subtract off the divergent parts (represented by $\mathcal{S}_{CT}$).

### v3

@@ C.2 Computation of $\gamma_{d-1}$-dependent part
The leading-order dependence of $\Sigma_{2}(k)$ on $k_{d}$ and $k_{d-1}$ can be extracted by setting $\mathbf{K}=0$. Hence, we will evaluate
@@
where $\delta_{q-k}=q_{d-1}-k_{d-1}+k_{d}^{2}+q_{d}^{2}-2\,k_{q}\,q_{d}$. Therefore,
$I_{3}$ $\simeq\int_{q}\frac{\delta_{q-k}}{\mathbf{Q}^{2}+\delta^{2}_{q-k}}\times\frac{1}{q_{d}^{2}}=\int\frac{d^{d-1}\mathbf{Q}\,dq_{d}\,du}{(2\,\pi)^{d+1}}\frac{u-k_{d-1}}{2\,\left[|\mathbf{Q}|^{2}+\left(u-k_{d-1}\right)^{2}\right]}\times\frac{2}{q_{d}^{2}}\quad(\text{where }u=\delta_{q}+2\,k_{d}\,q_{d}+k_{d}^{2})\,.$ (52)
Using the identity,
$\int_{-\infty}^{\infty}\frac{dw}{w^{2}+A}=\begin{cases}0&\text{ for }\mathcal{A}<0\\ \frac{\pi}{\sqrt{\mathcal{A}}}&\text{ for }\mathcal{A}>0\end{cases},$ (53)
for the Cauchy principal value of the integral, the $q_{d}$-integral is performed first to obtain
@@
$=\int\frac{d^{d-1}\mathbf{Q}\,d\tilde{u}}{(2\,\pi)^{d}}\frac{k_{d}^{2}+\tilde{u}-k_{d-1}}{\left[|\mathbf{Q}|^{2}+\left(k_{d}^{2}+\tilde{u}-k_{d-1}\right)^{2}\right]}\times\frac{\Theta(\tilde{u})}{2\,\sqrt{\tilde{u}}}\quad(\text{where }\tilde{u}=u-k_{d}^{2})$
@@
The factor $\Gamma\big(\frac{3}{2}-d\big)$ has a pole at $d=3/2$, which shows that this term has (1) a logarithmic divergence at $d=3/2$, and (2) a linear divergence at $d=5/2$, when we translate the divergences in the language of the Wilsonian cutoff $\Lambda\sim\mu$. We need to treat this result carefully by remembering that, in the dimensional regularization procedure, UV divergences of all degrees show up as the poles of $\Gamma$-functions. The degree of divergence can be understood by using an explicit UV cut-off in the language of the Wilsonian RG, which is denoted here by $\Lambda$. Although this term will play an important role for analysing UV-stable fixed-points, it should be discarded here, because we are considering the RG flows in the IR and this term represents an IR-irrelevant operator for the theory in $d=5/2-\epsilon$.
@@ Appendix D Renormalization Group flows under minimal subtraction scheme
Eq. (II) is supposed to be the physical action, defined at an energy scale $\mu\sim\Lambda$, consisting of the fundamental Lagrangian with non-divergent quantities. However, we have seen that the loop integrals lead to divergent terms and, in order to cure it, we employ the renormalization procedure, using dimensional regularization as the regularization method. In our dimensional regularization formalism, the UV-divergent terms are the ones arising in the $\epsilon\rightarrow 0$ limit. We use the minimal subtraction ($\rm MS$}) renormalization scheme to control the UV divergences [thooft, weinberg], which involves cancelling the divergent parts of the loop-contributions via adding appropriate counterterms. More precisely, we adopt the modified minimal subtraction ($\overline{\rm MS}$) scheme where, in addition to the divergent term, we absorb the universal term proportional to $\epsilon^{0}$ (that always accompanies the term with the $1/\epsilon$ pole) into the corresponding counterterm.
The action, consisting of the counterterms to absorb the singular terms, takes the following form:
@@
The counterterm-factors are given by the power series,
@@
Subtracting $\mathcal{S}_{CT}$ from the so-called bare action, $S_{\text{bare}}$, we obtain the renormalized action, which is the physical effective action of the theory re-written in terms of non-divergent quantum parameters. While the bare parameters can be divergent, the physical observables are the renormalized coupling constants, which are determined by the RG equations. The RG flows thus describe the evolution of the bare couplings as functions of the floating energy scale, $\mu\,g^{-l}$ (i.e., with respect to an increasing logarithmic length scale, $l$). To achieve this objective, we first define the bare (or fundamental) action
@@
consisting of the bare quantities, where the superscript “BB” has been used to denote the bare fields, couplings, frequency, and momenta. We now relate the bare quantities to the so-called renormalized quantities (without the superscript “BB”) via the multiplicative $Z_{\zeta}$-factors, such that
@@
Observing that there exists a freedom to change the renormalization of the fields and the renormalization of momenta without affecting the action, we have exploited it by requiring $k^{B}_{d}=k_{d}$. It is equivalent to measuring the scaling dimensions of all the other quantities relative to the scaling dimension of $k_{d}$. $\mathcal{S}$ now represents the renormalized action (also known as the Wilsonian effective action) because it consists of the renormalized quantities. Basically, we have written the fundamental action of our theory in two different ways [srednicki], which allows us to dump off the divergent parts in $\mathcal{S}_{CT}$.